Hello, I need help solving this problem. The data below represents the total compensation for 10...

Hello,

I need help solving this problem.

The data below represents the total compensation for 10 randomly selected CEOs and their​ company's stock performance. Analysis reveals a correlation coefficient of r = −0.2164. What would be the predicted stock return for a company whose CEO made​ $15 million? What would be the predicted stock return for a company whose CEO made​ $25 million?

Compensation ($ millions) Stock Return (%)
26.16 5.76
12.04 30.44
19.44 32.18
13.45 80.22
12.67 -8.72
12.27 2.64
26.01 4.81
14.64 11.03
17.52 4.18
14.49 11.57
Answer #1

and   

correlation cofficient (r) =

regression equation of Y on X

Y = -1.02X + 34.64

when CEO's compensation (X)= 15, then stock return (Y) will be

Y = -1.02(15) + 34.64 = 19.34 %

when CEO's compensation (X)= 25, then stock return (Y) will be

Y = -1.02(25) + 34.64 = 9.14 %